OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Part 2: Provide the solution to the system. 5x – y = –2 10x – 3y = –7

OpenStudy (kira_yamato):

5x - y = -2 -------(1) 10x - 3y = -7 -----(2) We will first make y the subject of the equation (1): y = 5x + 2 -------(3) Substituting (3) into (2): 10x - 3(5x + 2) = -7 10x - 15x - 6 = -7 -5x = -1 x = -1/5 Substitute x = -1/5 into (3) y = 5(-1/5) + 2 y = 1 Hence, the solution is x = -1/5, y =1.